SPEED DEPENDENT LINE SHAPES IN 1.61 μm AND 2.07 μm CO_{2} ATMOSPHERIC RETRIEVALS FOR THE OCO-2 MISSION

DAVID R. THOMPSON, LINDA R. BROWN, DAVID CRISP, YIBO JIANG, FABIANO OYAFUSO, \underline{KEEYOON SUNG}, CHARLES E. MILLER, VIJAY NATRAJ, DEBRA WUNCH, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr.,Pasadena, CA 91109, U.S.A.; D. CHRIS BENNER, V. MALATHY DEVI, The College of William and Mary, Williamsburg, VA 23187, U.S.A..

We are validating line parameters for CO_{2} at 1.61 μm and 2.07 μm using high resolution atmospheric spectra and a new retrieval algorithm [1, 2] being developed for the Orbiting Carbon Observatory (OCO-2) in order to estimate column-averaged mixing ratio of CO_{2}, X_{CO2}, to a sub-1% precision. This requirement demands highly accurate molecular line shape models. We tested a combination of line mixing [3] with speed dependent Voigt shapes [4,5] obtained from laboratory spectra using a state of the art multi-spectrum fitting procedure [6, 7]. The atmospheric tests were made with a diverse set of over 400 soundings including upward- and downward-looking FT-IR data from the Total Carbon Column Observing Network (TCCON), and the data from TANSO-FTS spectrometer aboard the Greenhouse gases Observing SATellite (GOSAT), respectively. The new absorption cross sections significantly reduced residuals in the spectral fit in the 2.07 μm region, while the effects on the 1.61 μm band are less definitive but still suggest some improvement. Overall these tests favor the adoption of the new models.^{a}^{b}

a[1] O’Dell C.W., et al. AMT 2012; 5:99-121. [2] Crisp, D., et al. AMTD 2012; 5:1 - 60. [3] Hartmann, J. M., et al. ACP 9:7303-7312. [4] Devi et al., J. Mol.Spec. 2007; 245:52-80. [5] Benner et al. 66th International Symposium on Molecular Spectroscopy, Columbus OH (2011). [6] Benner, D.C., et al., JQSRT 1995; 53(6):705 - 721. [7] Letchworth, K.L., et al. JQSRT, 107: 173 - 192.

bThe research described in this paper was performed at the Jet Propulsion Laboratory, California Institute of Technology and at The College of William and Mary under contracts with National Aeronautics and Space Administration. US Government Support Acknowledged.